express the following ratio having the same units in the numerator and denominator: 1 1/2 hours to 50 minutes. Please help
Answer by mathdoc314(58)

1 1/2 hours is an amount of time
50 minutes is another amount of time
These are both amounts of time.
You probably recognize that 1 1/2 hours is longer than 50 minutes, and 50 minutes is less than an hour. In fact 1 hour is exactly 60 minutes. Also half an hour is 30 minutes, and 1 1/2 hours is 90 minutes (one and a half hours is what that says).
All you gotta do here is convert hours to minutes and then re-express the ratio
90 minutes
50 minutes
90/50 = 9/5 = 1 4/5 (nine fifths, or one and four fifths).
that is the ratio.
The first amount is 9/5 or 1 4/5 as big as the second amount.